OpenStudy (anonymous):

Solve the following equation algebraically: sin²x - √2cosx = cos²x + √2cosx + 2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

cos x=t cos^2x =t^2 sin^2x=1-cos^2x=1-t^2 1-t^2-sqrt(2)t=t^2+sqrt(2)t+2 2t^2+2sqrt(2) t-1=0

OpenStudy (anonymous):

\[t=(-2\sqrt{2}\pm \sqrt{8+8})/4\]

OpenStudy (anonymous):

\[t= (-2\sqrt{2}\pm4)/4\]

OpenStudy (anonymous):

\[x=\cos ^{-1}[(-2\sqrt{2}\pm 4)/4]\]